# Stochastic gradient descent-based inference for dynamic network models with attractors

### v2

@@ 4 Simulation studies
We also compare the computational efficiency of the proposed SGD algorithm with MCMC algorithm. When both run on the same CPU resources, the MCMC algorithm took about an hour to obtain 50,000 samples when the number of nodes $n=100$ while the proposed SGD-based algorithm only took fewer than 5 minutes. In the X data analysis in the following section, on the reduced data set with 200 nodes, the MCMC took about four hours (CPU) while the proposed SGD-based algorithm only took fewer than 10 minutes to run(GPU). In addition, the proposed SGD-based algorithm took about 20 minutes when $n=1000$ as well (GPU).
@@ 5.1 Exploratory Analysis
5.2 Implementation Details
We use gradient descent with momentum and choose different learning hyperparameters for the latent position parameters and the global parameters. With gradient descent, we use all the terms in the log posterior function instead of taking a sampled posterior function. We stop running gradient descent when the parameter updates drop below a threshold of $\epsilon=10^{-4}$, a condition we check every 100 gradient steps. Priors for $\alpha$ and $\delta$ were chosen to be $\mathcal{N}(0,100)$ to keep it flat and uninformative. We chose the priors for $\gamma^{w}$ and $\gamma^{b}$ to be $\mathcal{N}(0.5,100)$ and $\mathcal{N}(-0.5,100)$ to reflect the prior belief of polarization, however these are also quite uninformative given the large variance. We fix $\tau^{2}$ at 10, $\sigma^{2}$ at 1 and $\phi^{2}$ at 10.
@@ 5.3 Results
Figure 4: The point estimates of the latent positions with the extend model. Blue-Democrats, Red-Republicans. At each time, two parties consistently occupy different halves of the space. The Democrats were flocking during the whole time. Conversely, the Republicans initially displayed a similar flocking behavior, but gradually began to disperse around year 2016.
@@
We fit our model with time-invariant parameters to the full X data set from year 2010 to 2020. The point estimates and the posterior standard deviation of model parameters are provided in Table 2. Point estimates were produced using the SGD methodology of Section 3.1 and standard deviation estimates were produced using the methodology of Section 3.2. In the supplementary file, we present two diagnostic tests demonstrating that the underlying assumptions of the variance estimation method proposed in Section 3.2 are largely accurate for this dataset, i.e., the diagnostic tests do not show evidence of violating the underlying assumptions of the variance estimation method.
The point estimates of persistence parameter is 1.08, implying that there is a higher likelihood for it to also appear at time $t$ if an edge is present at time $t-1$. The between-group attraction is -0.22, indicating that there is polarization between the two parties. The within group coefficient is 0.34 for Democrats, and -0.11 for Republicans, indicating that the Democrats were flocking, while the Republicans were polarizing within their own group. Importantly, we note that in the original model in Zhu et al. (2023), only members that are present at all times are kept. The parameter estimation for attractions within group indicated that while both parties have moved away from one another, they generally flocked to their own. By incorporating all members into the extended model, we effectively eliminate the potential for selection bias, ensuring a more accurate and comprehensive analysis of the data. As a result of this approach, we have uncovered that the force within the Republican group is negative, in contrast to Zhu et al. (2023).
@@
Figure 5: The trajectory of the mean of the latent positions of the members in each party. The lighter color are results obtained from the reduced data set and the darker color are results obtained from the full data set. In both cases, the average position of the two parties gradually and consistently approached each other from 2011 to 2016 and after 2016, both of them took a U-turn and moved apart.
@@
To analyze change in attraction, repulsion, we fit a series of models that allow a change-point to vary from 2011 to 2019. The resulting fitted models with different change-points were compared based on Bayesian Information Criterion (BIC). The one with the lowest BIC value was selected. We identified 2012 as the year of change-point (See supplementary file for more details on this). Figure 7 illustrates the evolution of within-group attraction and repulsion for Democrats and Republicans, and between-group attraction and repulsion. The within-Republican coefficient is positive in the first time period and is negative in the second time periods from 2012 to 2020. This suggests polarization within Republican members started to rise within the period 2012-2020. The within-Democrat coefficient is positive in both time periods, although its magnitude decreases a bit in the second time period from 2012 to 2020. The between-group coefficient (orange bars) is not statistically significant at the 95% confidence level from 2010 to 2012 and is negative in the second time periods from 2012 to 2020.
@@ Acknowledgement
This supplementary document contains two sections. In Section 7 we present the proofs of the theoretical results that appear in the main body of the paper. In Section 8 we present some additional statistical tests and diagnostics to back up the validity of the numerical results reported in the main body of the paper.
7 Proofs
@@
Consider the multivariate Gaussian distribution setting presented above, and let $x_{2}$ be scalar. Then, for some fixed $\tilde{x}_{2}\neq\mu_{2}$.
[…]
@@ 8.2.2 Linearity test
Figure 9: We choose $\theta_{1}$ to be $\gamma^{w}_{R}$, and set $\eta$ to be -0.03, -0.02, -0.01, 0.01, 0.02, 0.03. We calculate $\frac{\mu(X_{j}|X_{1}=\mu_{1}+\eta)-\mu_{j}}{\eta}$ for each latent position parameters under each $\eta$. The latent position parameters are ordered by the mean of its slopes and plot the mean maximum and minimum of its slopes (left) and the slopes of 95% of the latent position parameters after removing the tail 5% that have a very large or very small slopes.
@@
We use the results of the previous test, set $\theta_{1}$ to be $\gamma^{w}_{R}$, and $\eta$ to take values in the set $\{-0.03,-0.02,-0.01,0.01,0.02,0.03\}$. We calculate $\frac{\mu(X_{j}|X_{1}=\mu_{1}+\eta)-\mu_{j}}{\eta}$ for each latent position parameters and each $\eta$. In Figure 9 we order the latent position parameters by the mean of its slopes and plot the mean maximum and minimum of its slopes (left) and the slopes of 95% of the latent position parameters after removing the tail 5% that have a very large or very small slopes. We can see that the slopes, or the correlated changes for each latent position parameters when we change $\gamma^{w}_{R}$ are close under different choice of $\eta$. Therefore the diagnostic plot does not show substantial evidence of violating the underlying assumptions of the variance estimation method.

### v3

@@ 4 Simulation studies
As a comparison of the computational efficiency of the proposed SGD algorithm with MCMC algorithm, in the X data analysis in the following section, on the reduced data set with 200 nodes, the MCMC took about four hours (CPU) while the proposed SGD-based algorithm only took fewer than 10 minutes to run (GPU). The supplementary file includes a more thorough comparison among MCMC and SGD-based estimation (CPU and GPU).
In the supplementary file, we also include a simulation study to demonstrate the algorithm’s robustness in the scenario where nodes are allowed to leave the network over time, i.e., in the case where the vertex set $V_{t}$ changes dynamically in time.
@@ 5.2 Results
Figure 5: The point estimates of the latent positions with the extended model. The dots that mainly occupy the left portion of each plot represent Democrats and the dots that mainly occupy the right portion of each plot represent Republicans. At each time point, the two parties consistently occupy different halves of the space. The Democrats exhibited clustering behavior throughout the period. Conversely, the Republicans initially displayed a similar flocking behavior, but gradually began to disperse around the year 2016.
@@
We fit our model with time-invariant parameters to the full X data set from year 2010 to 2020. The AUC values computed at each year are all above 0.974, and the overall AUC value computed across all times is 0.985, providing evidence that our model fits the data very well. The point estimates and the posterior standard deviation of model parameters are provided in Table 2. Point estimates were produced using the SGD methodology of Section 3.1 and standard deviation estimates were produced using the methodology of Section 3.2. In the supplementary file, we present two diagnostic tests demonstrating that the underlying assumptions of the variance estimation method proposed in Section 3.2 are largely accurate for this dataset, i.e., the diagnostic tests do not show evidence of violating the underlying assumptions of the variance estimation method.
The point estimates of persistence parameter is 1.08, implying that there is a higher likelihood for it to also appear at time $t$ if an edge is present at time $t-1$. The between-group attraction coefficient $\gamma^{b}$ is -0.22, indicating that there is polarization between the two parties. The within group coefficients $\gamma^{w}$ are 0.34 for Democrats, and -0.11 for Republicans, indicating that the Democrats were flocking, while the Republicans were polarizing within their own group. Importantly, we note that in the original model in Zhu et al. (2023), only members that are present at all times are kept. The parameter estimation for attractions within group indicated that while both parties have moved away from one another, they generally flocked to their own. By incorporating all members into the extended model, we effectively eliminate the potential for selection bias, ensuring a more accurate and comprehensive analysis of the data. As a result of this approach, we have uncovered that the force within the Republican group is negative, in contrast to Zhu et al. (2023).
@@
Figure 6: The trajectory of the mean of the latent positions of the members in each party. The average position of the two parties gradually and consistently approached each other from 2011 to 2016 and after 2016, both of them took a U-turn and moved apart.
@@
To analyze the changes in attraction, repulsion, we fit a series of models that allow a change-point to vary from 2011 to 2019. The resulting fitted models with different change-points were compared based on Bayesian Information Criterion (BIC). The one with the lowest BIC value was selected. We identified 2012 as the year of change-point (See supplementary file for more details on this). Figure 8 illustrates the evolution of within-group attraction and repulsion for Democrats and Republicans, and between-group attraction and repulsion. The within-Republican coefficient is positive in the first time period and is negative in the second time periods from 2012 to 2020. This suggests polarization within Republican members started to rise within the period 2012-2020. The within-Democrat coefficient is positive in both time periods, although its magnitude decreases a bit in the second time period from 2012 to 2020. The between-group coefficient (orange bars) is not statistically significant at the 95% confidence level from 2010 to 2012 and is negative in the second time periods from 2012 to 2020.
The observed patterns of increasing polarization between the Democratic and Republican parties align with findings from previous studies. For instance, Andris et al. (2015) documented similar trends through the analysis of co-sponsorship patterns in Congress, while Garimella and Weber (2017) identified heightened polarization in the retweet networks of U.S. politicians on Twitter. These studies, along with our findings, highlight political polarization in different environments in U.S..
@@ 6 Discussion
With the increased efficiency of the proposed method, extending the model to accommodate more than two groups with attractors becomes more feasible. This extension is particularly useful in contexts such as multi-party political systems in Germany and Canada. For a small number of groups, the computational impact is expected to be minimal. Moreover, our library has been developed with this flexibility in mind and allows for such an extension.
@@ Acknowledgement
Supplementary Materials
Appendix:
Contains proofs of the theoretical results, additional statistical tests and diagnostics to support the numerical results, implementation details of the algorithm, performance/efficiency studies of the algorithm, and a simulation study demonstrating the algorithm’s robustness against node turnover. (supp.pdf, PDF)
Source Code:
The source code is available in the GitHub repository. The current library implementation is specifically designed for a 2-party system.
(https://github.com/KolaczykResearch/SGD4CLSNA, GitHub repository)
@@
This supplementary document is organized as follows. In Section 7 we present the proofs of the theoretical results that appear in the main body of the paper. In Section 8 we present some additional statistical tests and diagnostics to back up the validity of the numerical results reported in the main body of the paper. In Section 9, we provide implementation details for the algorithm. In Section 10, we compare computational times among MCMC and the newly proposed SGD-based algorithm (in both CPU and GPU settings). We also comment upon the impact of hyper-parameters choices on convergence times. In Section 11 we present a simulation study to illustrate the algorithm’s robustness when nodes exit the network at different turnaround levels.
7 Some Useful Results of Multivariate Normal Distribution
More specifically, the covariance of the conditional distribution $\text{Cov}(x_{1}|x_{2})$ is a constant as a function of $x_{2}$, indicating that the shape of the conditional distribution is independent of the specific value of $x_{2}$. Additionally, writing $p(x_{1}|x_{2})$ as the p.d.f of the conditional distribution of $x_{1}|x_{2}$, the mean $\mu_{1|2}=\mu_{1}+\Sigma_{12}\Sigma_{22}^{-1}(x_{2}-\mu_{2})$ is the median, and mode of the distribution. The maximum of the p.d.f. function $p(x_{1}=\mu_{1|2}|x_{2})$ is a constant as a function of $x_{2}$.
@@ 7 Some Useful Results of Multivariate Normal Distribution
In particular, the marginal distribution is proportional to the joint distribution evaluated on the curve of the conditional mean of $x_{1}$ given $x_{2}$.
@@
Theorem 7.1 provides an expression for the shape of the marginal distribution from the joint distribution, when the joint distribution is a multivariate normal distribution.
Next, we proceed with an important result demonstrating that for a normal distribution the ideas of Theorem 7.1 yield a formula for the variance of $x_{2}$ that we can then be turned into a practical SGD-based algorithm for uncertainty quantification. In particular, recall that in a univariate normal distribution, two key parameters - the mean and variance - describe the distribution fully. By considering two distinct points from the distribution, we can deduce these parameters.
@@
$\text{Var}(x_{2})=\frac{1}{2}\frac{(\mu_{2}-\tilde{x}_{2})^{2}}{\log p(x_{1}=\mu_{1},x_{2}=\mu_{2})-\log p(x_{1}=\mu_{1|2},x_{2}=\tilde{x}_{2})}$ (17)
@@
Corollary 7.1.2.
$\Sigma_{12}=\frac{\Sigma_{22}}{(\tilde{x}_{2}-\mu_{2})}(\mu_{1|2}-\mu_{1}),$
The methodology established for a generic normal distribution $p(x_{1},x_{2})$ is adapted to estimate variance in the posterior distribution $\pi(\cdot|Y)$, effectively applying the same theoretical concepts to the Bayesian analysis under a Laplace approximation.
@@ 8.2.2 Linearity test
Figure 10: We choose $\theta_{1}$ to be $\gamma^{w}_{R}$, and set $\eta$ to be -0.03, -0.02, -0.01, 0.01, 0.02, 0.03. We calculate $\frac{\mu(X_{j}|X_{1}=\mu_{1}+\eta)-\mu_{j}}{\eta}$ for each latent position parameters under each $\eta$. The latent position parameters are ordered by the mean of its slopes and plot the mean maximum and minimum of its slopes (left) and the the slopes of 95% of the latent position parameters after removing the tail 5% that have a very large or very small slopes.
@@
We use the results of the previous test, set $\theta_{1}$ to be $\gamma^{w}_{R}$, and $\eta$ to take values in the set $\{-0.03,-0.02,-0.01,0.01,0.02,0.03\}$. We calculate $\frac{\mu(X_{j}|X_{1}=\mu_{1}+\eta)-\mu_{j}}{\eta}$ for each latent position parameters and each $\eta$. In Figure 10 we order the latent position parameters by the mean of its slopes and plot the mean maximum and minimum of its slopes (left) and the the slopes of 95% of the latent position parameters after removing the tail 5% that have a very large or very small slopes. We can see that the slopes, or the correlated changes for each latent position parameters when we change $\gamma^{w}_{R}$ are close under different choice of $\eta$. Therefore the diagnostic plot does not show substantial evidence of violating the underlying assumptions of the variance estimation method.
9 Implementation Details
We use gradient descent with momentum and choose different learning hyperparameters for the latent position parameters and the global parameters. With gradient descent, we use all the terms in the log posterior function instead of taking a sampled posterior function. We stop running gradient descent when the parameter updates drop below a threshold of $\epsilon=10^{-4}$, a condition we check every 100 gradient steps. Priors for $\alpha$ and $\delta$ were chosen to be $\mathcal{N}(0,100)$ to keep it flat and uninformative. We chose the priors for $\gamma^{w}$ and $\gamma^{b}$ to be $\mathcal{N}(0.5,100)$ and $\mathcal{N}(-0.5,100)$ to reflect the prior belief of polarization, however these are also quite uninformative given the large variance. We fix $\tau^{2}$ at 10, $\sigma^{2}$ at 1 and $\phi^{2}$ at 10. We choose $p=2$, following Zhu et al. (2023). The choice of two dimensions aligns with DW-NOMINATE, a widely recognized model of congressional ideology. This model demonstrates that two dimensions can account for up to 90% of the variation in roll call voting (McCarty et al., 2016).
For the step size $\lambda$, we started by conducting a search to identify the optimal initial step size for our model by observing the loss curves for various candidates. The goal was to identify a step size that avoided slow convergence and prevented the loss from exploding. Once the best initial step size was determined, we proceeded with training the model. During training, if the loss plateaued, we halved the step size to promote further reduction in the loss. After one such adjustment, further halving did not yield significant improvements, so we concluded the training process at that point. To choose $\eta_{\alpha}$, we take $\alpha$ as an example. We start by selecting an initial guess for $\eta_{\alpha}$ such that $\log\pi(\alpha^{*})-\log\pi(\alpha^{*}+\eta_{\alpha})$ is neither too large nor too small, aiming for a value between 10 and 50. Here, $\log\pi(\alpha^{*})$ denotes the log probability density function $\pi$ evaluated at the mode with $\alpha$ set to be $\alpha^{*}$, while $\log\pi(\alpha^{*}+\eta_{\alpha})$ represents the same function evaluated at the mode with $\alpha$ set to be $\alpha^{*}+\eta_{\alpha}$. This range ensures that the perturbation is sizable enough to influence the parameters meaningfully but not so large that they deviate excessively from the optimum. In our case, we opted for a uniform $\eta_{\alpha}$ for all parameters for which we wanted to estimate variance, and this approach worked well.
10 Algorithm Comparison and Analysis
Figure 11: Comparison of Convergence Times for SGD (CPU and GPU) and MCMC Algorithms Across Different Network Sizes.
Figure 11 illustrates the convergence times for SGD-based algorithms executed on a CPU and an Nvidia GPU A40, alongside the convergence times for the MCMC algorithm. The x-axis denotes the number of nodes in the network, while the y-axis represents the convergence time in minutes. The blue curve corresponds to the convergence times of the SGD algorithm when executed on a CPU, the red curve represents the convergence times of the same algorithm when executed on a GPU, and the green curve shows the convergence times for the MCMC algorithm. The SGD algorithm provides approximately a 30-fold reduction in convergence time compared to the MCMC algorithm. Additionally, parallelization further improves the scalability of the SGD method: executing the SGD algorithm on the Nvidia GPU A40 yields substantial efficiency gains, resulting in around a 20-fold reduction in convergence time compared to its execution on the CPU. Note that the reported convergence times for the SGD algorithm include both the point estimation step and the variance estimation step.
10.1 Impact of Hyperparameters on Convergence Time
It is important to understand how hyperparameters, such as step size ($\lambda$), affect the convergence speed. Choosing the correct step size ($\lambda$) and momentum in SGD is crucial, as it significantly affects the convergence behavior. Specifically, if the step size is too large, the algorithm may diverge, causing the loss to explode. If the step size is too small, the algorithm will converge very slowly. To address this, we fixed the momentum parameter at 0.99 and focused on tuning the step size. For the step size $\lambda$, we started by conducting an empirical search to identify an optimal initial value. We observed the loss curves for various step size candidates, aiming to find a step size that avoids both slow convergence and loss explosion. After determining the best initial step size, we began training the model. During training, if we noticed that the loss plateaued, we halved the step size. This halving process was repeated until the log posterior function convergences. Once the correct magnitude is established, fine-tuning the step size only results in marginal gains in convergence speed. For the computational time plot presented above, we used the step size of $0.001$ for all latent position parameters and $0.001$ for all global parameters, while using the sign of the gradient (+1/-1) for updating the global parameters instead of the actual gradient values.
10.2 Computational Complexity and Factors Contributing to Speedup
Each iteration of the proposed SGD-based algorithm has a computational complexity of $O(T\times N^{2})$ for both time and space, where $T$ is the number of time points in the network time-series and $N$ represents the number of nodes in each network. This complexity comes from the requirement to process all node pairs across each time point. Each iteration of the MCMC algorithm also has the same computational complexity of $O(T\times N^{2})$. The significant speedup of the SGD-based algorithm predominantly stems from two key factors:
  1. 1.
    Parallelization on GPU: The ability to leverage the parallel processing capabilities of GPUs significantly reduces the time spent per iteration. This factor provides around a 20-fold increase in performance compared to the CPU execution.
  2. 2.
    Fewer Iterations to Converge: The SGD-based algorithm requires far fewer iterations to achieve convergence compared to the MCMC algorithm. This results in approximately a 30-fold increase in performance.
11 Varying percentage leaving
$\alpha=1$ $\delta=2$ $\gamma^{w}=0.25$ $\gamma^{b}=0.5$
Turnover=0% 0.94 (0.004) 0.005 ($<$0.001) 1.98 (0.005) 0.005 ($<$0.001) 0.24 (0.025) 0.028 (0.007) 0.50 (0.025) 0.026 (0.008)
Turnover=10% 0.94 (0.006) 0.005 ($<$0.001) 1.98 (0.007) 0.005 ($<$0.001) 0.24 (0.032) 0.030 (0.007) 0.50 (0.030) 0.027 (0.008)
Turnover=20% 0.94 (0.006) 0.005 ($<$0.001) 1.98 (0.007) 0.006 ($<$0.001) 0.25 (0.041) 0.032 (0.007) 0.49 (0.034) 0.028 (0.008)
Turnover=30% 0.94 (0.008) 0.005 ($<$0.001) 1.98 (0.005) 0.007 ($<$0.001) 0.24 (0.044) 0.034 (0.008) 0.50 (0.034) 0.030 (0.010)
Turnover=40% 0.94 (0.007) 0.005 ($<$0.001) 1.98 (0.008) 0.008 ($<$0.001) 0.26 (0.042) 0.036 (0.008) 0.48 (0.029) 0.030 (0.009)
Turnover=50% 0.94 (0.008) 0.005 ($<$0.001) 1.98 (0.008) 0.009 ($<$0.001) 0.25 (0.042) 0.038 (0.008) 0.49 (0.031) 0.032 (0.010)
Turnover=60% 0.94 (0.007) 0.005 ($<$0.001) 1.98 (0.010) 0.011 ($<$0.001) 0.27 (0.036) 0.039 (0.008) 0.48 (0.027) 0.031 (0.009)
Turnover=70% 0.94 (0.007) 0.005 ($<$0.001) 1.98 (0.014) 0.015 ($<$0.001) 0.26 (0.030) 0.043 (0.008) 0.48 (0.025) 0.033 (0.011)
Turnover=80% 0.95 (0.005) 0.005 ($<$0.001) 1.99 (0.024) 0.022 ($<$0.001) 0.26 (0.056) 0.049 (0.008) 0.49 (0.038) 0.034 (0.012)
Table 4: Posterior-based mean (empirical standard deviation) for point estimation and variance estimation for parameters $\alpha$, $\delta$, $\gamma^{w}$, and $\gamma^{b}$, based on $n=500$ nodes with $T=10$ time points. The table presents the robustness of the estimation method across varying proportions of node turnover, ranging from 0% to 80%. Despite increasing turnover, the point estimates remain consistent. The variance of the estimates for $\delta$, $\gamma^{w}$, and $\gamma^{b}$ increases due to the reduction in sample size, which is also captured by the variance estimates appropriately.
Table 4 presents the posterior-based mean (empirical standard deviation) for point estimation and variance estimation, based on a CLSNA model with $n=500$ nodes, similar to the size of a US Congress, across varying proportions of nodes entering and leaving the network to demonstrate the robustness of the method. Similar to Congress, at each time step, a random subset of nodes are dropped out, while an equal number of new nodes are randomly initialized according to the model’s prior distribution. The results indicate that both the point estimates and variance estimates are robust across different turnover levels. As the turnover proportion increases, the variance of the estimates for $\delta$, $\gamma^{w}$, and $\gamma^{b}$ also increases, which can be attributed to the smaller sample sizes available for estimating these parameters. This shows the method’s capability to provide robust point estimates with varying proportions of nodes entering and leaving while effectively capturing the increase in uncertainty as the turnover proportion rises.